Atmospheric pressure at sea level is 101 kPa and the density of sea water is 1020 kgm-3, therefore the total pressure 42 m underwater is:

A 420 kPa

B 521 kPa

C 4327 MPa

D 144 kPa


1
Expert's answer
2022-03-24T11:00:36-0400

The total pressure 42 m underwater is

"p=p_0+g\rho h\\
=101*10^3+9.8*1020*42\\
=5.21*10^5\:\rm Pa"

Answer: B